Use Scenario 1 to answer questions 1A, 1B, and 1C below (2 - 4).

Scenario 1: Newmont Mining Company is considering investing in a piece of land that may be purchased for $610,000 (today) to be open pit-mined for copper for the next 10 years. The mine is expected to produce income of $200,000 at the end of each year for 10 years. At the end of 10 years, the land will be restored in line with federal law requirements , which will cost $175,000 at that time. Assume Newmont Mining Company’s MARR is 10% per year, compounded annually.

1A. Calculate the annual worth for this investment.

1B. Given the following information, calculate the internal rate of return for this investment using interpolation.

i = 28%, NVP = 28960.28

i = ?, NPV = ?

i = 30%, NPV = -4386.28

1C. Is the land investment acceptable?
Answer #1

Initial Cost=C=$610,000

Annual Income=R=$200,000 per year

Restoration Cost=S=$175,000

Time period=n=10 years

MARR=i=10%=0.10

1 A)

AW=-C*(A/P,0.10,10)+R-S*(A/F,0.10,10)

AW=-610000*(A/P,0.10,10)+200000-175000*(A/F,0.10,10)

Let us calculate interest factors

(A/P,1,n) = 1 -

0.10 (A/P,0.10, 10) = 1 = 0.162745 (1+0.10)10

(A/F, i, n) = (1 + i)n-1

0.10 (A/F,0.10, 10) = (1 + 0.10) 10_= 0.062745

AW=-610000*0.162745+200000-175000*0.062745=$89745.18 or say $89745

1B)

iL=28%

NPVL=28960,28

iH=30%

NPVH=-4386.28

NPV at IRR=0

IRR=i=?

Interpolation gives us

IRR=iL+(NPVL*(iH-iL)/(NPVL-NPVH))

IRR=28%+(28960*(30%-28%)/(28960.28+4386.28))=29.74%

IC)

In this case

AW is positive at MARR,

IRR is higher than MARR

So, we can say that investment is acceptable.